OpenStudy (anonymous):

how do i solve this equation: x/2+x/3=5

OpenStudy (anonymous):

multiple 6 and find x out.

OpenStudy (anonymous):

Find the common denominator for x/2 and x/3. Then add them together. Once you get the sum, multiply both sides of the equation by the denominator and solve. Do you understand?

OpenStudy (anonymous):

yeah,i"ll try working it out now

OpenStudy (anonymous):

OK - let me know if you get stuck

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

\[x/2+x/3= 3/6+2/6= 5/6\]

OpenStudy (anonymous):

is that right so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5x/6

OpenStudy (anonymous):

because it was 3x/6+2x/6

OpenStudy (anonymous):

okay,after that im confused on what to do next.

OpenStudy (anonymous):

so you have 5x/6=5. You need to eliminate the 6 in the denominator. Since you are dividing by 6, perform the inverse operation and multiply both sides by 6.

OpenStudy (anonymous):

so it would be 30 divided by 5x ?

OpenStudy (anonymous):

Close. Multiplying the left side by 6 leaves you with 5x. Multiplying the right side by 6 gives you 30. So now you have 5x=30. I bet you can solve that.

OpenStudy (anonymous):

so do i divide each side by 5 now?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

okay and that should get me the answer right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay thanks! :)

OpenStudy (anonymous):

you might want to go back and review the steps for solving this problem - you'll probably see some more similar to this one. Good luck!